OpenStudy (anonymous):

This circle is centered at the point (4, 5), and the length of its radius is 3. What is the equation of the circle

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x - h)^2 + (y - k)^ 2 = r^3 }\)

Parth (parthkohli):

r is the radius and (h,k) is the vertex.

Parth (parthkohli):

Just substitute the values.

OpenStudy (anonymous):

\[(x-4)+(y-5)=9\]

OpenStudy (anonymous):

thanks you guys yall are a great help